I'm not sure how to do this problem, can you please help?
Thanks!
Consider:
x + 6x^(1/2) + 9 = 0
Can we use the quadratic formula to solve this equation, and if so how?

Answer by Nate(3500)

y = x^(1/2) or sqrt(x)
(x^(1/2))^2 + 6x^(1/2) + 9 = 0
(y)^2 + 6y + 9 = 0
(y + 3)(y + 3) = 0
(sqrt(x) + 3)^2 = 0
sqrt(x) + 3 = 0
sqrt(x) = -3
x = 9
Plug:
9 + 6*sqrt(9) + 9 = 0
18 +- 18 = 0
When the sqrt(9) is negative, the answer is correct. Otherwise, no solution.